Let ABCDEF be a regular hexagon.If AD=xBC and CF=yAB, then xy= ?

Grade:12

1 Answers

Arun
5 years ago
given AB =a

BC =b   here a and b are vector

EF=-b

DE=-a

and we know that AD=2BC

=2b

AD=AB+BC+CD

2b =a+b+CD

CD =b-a

and FA=-CD

=a-b

Ac=AB+BC

=a+b

AF=CD=b-a

and CF=2DE

=-2a